OpenStudy (anonymous):

A movie and an order of popcorn cost $11.25 together. What is the price of the movie if it costs twice as much as the popcorn? Answer $7.50 $3.75 $8.25 $4.00

OpenStudy (anonymous):

@dpgreen

OpenStudy (anonymous):

@iNeedTheHelp

OpenStudy (anonymous):

8:25

OpenStudy (anonymous):

$7.50

OpenStudy (anonymous):

two different answers???

OpenStudy (anonymous):

$7.50

OpenStudy (anonymous):

i will take that answer

OpenStudy (anonymous):

use the equation 11.25=2x+x which is the same as saying 11.25=3x and then divide 11.25 by 3 for 3.75time that by 2 to get the price of the movie

OpenStudy (anonymous):

m- cost of movie p- cost of popcorn m+p= 11.25 m= 2p 2p +p = 11.25 3p = 11.25 p= $3.75 m = 2p m= 2( 3.75) m= $7.50 the movie costs $7.50 so subtract that from the total $11.25 to get the cost of the popcorn which is $3.75